Let P(X 14) = 0.31, P(14 < X 27) = 0.49, and P(X> 22) = 0.46. (a) Find P(X > 14) (b) Find P(X S 27). (c) Find P(14 < X 22). (d) If P(18 < X 27) 0.47, find P(X 18). Round your answers to two decimal places (e.g. 98.76).
Answer #1

a) P(X>14) = 1-0.31 = 0.69

b) P(X<=27) = 0.49-0.31 = 0.18

c) P(14<X<=22) = (1-0.46)-0.31 = 0.23

d) P(X<=18) = 0.47-0.18 = 0.29